# Information-Theoretic Limits of Bistatic Integrated Sensing and Communication

### v3

@@ III-D Capacity-distortion region for a bistatic ISAC broadcast channel model
In Fig. 1, if the ComRx is statistically stronger than the SenRx, i.e., $Z$ is degraded with respect to $Y$, then the channel model degenerates to $p(y,z|x,s)=p(y|x,s)p(z|y)$, which is similar to the channel model of the degraded broadcast channel. Thus, we further derive that the capacity-distortion region of the bistatic ISAC broadcast systems, where the transmitter sends the message to the ComRx and the SenRx, the ComRx and the SenRx decode the information, and SenRx also needs to estimate the state of the channel. We propose a two-receiver state-dependent memoryless broadcast channel (SDMBC) model $\left(\mathcal{X},\mathcal{S},p\left(y,z|x,s\right),\mathcal{Y}\times\mathcal{Z}\right)$, which consists of four finite sets $\mathcal{X},\mathcal{S},\mathcal{Y},\mathcal{Z}$ and a collection of pmf $p\left(y,z|x,s\right)$ on $\mathcal{Y}\times\mathcal{Z}$. As shown in Fig. 3, the ISAC Tx sends a message $W=(W_{0},W_{1},W_{2})$, where $W_{0}$ is the common message to the ComRx and the SenRx, $W_{1}$ and $W_{2}$ are the private message to the ComRx and the SenRx, respectively.
@@ IV-A Example 1
We consider a binary channel model, where the signal received by the ComRx and the SenRx are given by $Z=Y=S\oplus X$ with binary alphabets $\mathcal{X}=\mathcal{S}=\mathcal{Y}=\{0,1\}$ and the state $S$ is Bernoulli$(q)$, for $q\in(0,1)$. The Hamming distortion measure $d(s,\hat{s})=s\oplus\hat{s}$ is employed to characterize the sensing accuracy.
In the following, we apply the theoretical results in section III to derive the capacity-distortion function. Let $\bar{e}=1-e$, $P(X=0)=p\in[0,1]$, and $e\ast p=e\bar{p}+\bar{e}p$. Besides, ${\rm{H}}_{2}(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$ is the binary entropy function and $\rm{H}_{2}^{-1}(\cdot)$ refers to the inverse function of the binary entropy function $\rm{H}_{2}(\cdot)$ with the variable taking the value in $[0,1/2]$. It is seen from the channel model in Example 1 that $Z$ is a degraded form of $Y$, i.e., the Markov chain $X-(Y,S)-Z$ holds. It follows from Corollary 3 that the capacity-distortion function is given by
$C(D)=\max_{P_{UX}}\big{\{}H(X|U)+I(U;Z)\big{|}\mathbb{E}\big{[}d(S,\hat{S}^{*}(U,Z))\leq D\big{]}\big{\}},$ (22)
@@
Without loss of generality, we assume that $q\geq 1/2$. Applying Theorem 3 to Example 1, we obtain that $R=I(X;Y|U,S)+I(U;Z)=H(X|U)+H(Z)-H(Z|U)$ and the distortion $\mathbb{E}\big{[}d(S,\hat{S}^{*}(U,Z))=\sum_{u,z}\min_{s}p(s,u,z)$. Similar to the proof of the capacity region of a binary symmetric broadcast channel (BS-BC), we get that $R={\rm{H}}_{2}(\alpha)+1-{\rm{H}}_{2}(\alpha\ast q)$ is achievable by applying the superposition coding and successive cancellation decoding, i.e., $X=U\oplus V$, where $U$ and $V$ are independent, $U$ is Bernoulli$(\beta)$, and $V$ is Bernoulli$(\alpha)$ for $\beta\in[0,1]$. Correspondingly, the distortion $D=\min\{\alpha\bar{q},\bar{\alpha}q\}+\min\{\bar{\alpha}\bar{q},\alpha q\}$.
In the following, we prove the converse. By leveraging the fact that there exists an $\alpha\in[0,1/2]$ such that $0\leq H(Z)-H(Z|U)\leq 1-{\rm{H}}_{2}(q)=1-{\rm{H}}_{2}(\alpha\ast q)$ and the Mrs. Gerber’s Lemma in [45], we get that ${\rm{H}}_{2}(\alpha\ast q)=H(Z|U)\geq{\rm{H}}_{2}\big{(}{\rm{H}}_{2}^{-1}\big{(}{\rm{H}}_{2}(X|U)\big{)}\ast q\big{)}$. According to the monotonicity of binary entropy function and combining the facts that $\alpha\ast q\geq 1/2$ and $\alpha\leq e$ if $\alpha\ast q\geq e\ast q$, we have $H(X|U)\leq{\rm{H}}_{2}(\alpha)$ for $\alpha\in[0,1/2]$, which leads to $R\leq{\rm{H}}_{2}(\alpha)+1-{\rm{H}}_{2}(\alpha\ast q)$. Furthermore, since $\left|\mathcal{U}\right|\leq|\mathcal{X}|+1=3$, we set $P(U=i)=\beta_{i},i=0,1,2$, where $\sum_{i=0}^{2}\beta_{i}=1$. Thus, we have $D\geq D_{\min}(\alpha_{0},\alpha_{1},\alpha_{2})=\sum_{i=0}^{2}\beta_{i}\big{(}\min\{\alpha_{i}\bar{q},\bar{\alpha_{i}}_{1}q\}+\min\{\bar{\alpha}_{i}\bar{q},\alpha_{i}q\}\big{)}$, where $p(x=0|u=i)=\alpha_{i},i=0,1,2$. From the above expression about $D_{\min}$, we further get that $D_{\min}(\alpha_{0},\alpha_{1},\alpha_{2})\geq\min_{\alpha_{i},i=0,1,2}\{\min\{\alpha_{i}\bar{q},\bar{\alpha_{i}}q\}+\min\{\bar{\alpha_{i}}\bar{q},\alpha_{i}q\}\}$. Therefore, there exists an $\alpha\in[0,1/2]$ such that $D_{\min}(\alpha_{0},\alpha_{1},\alpha_{2})\geq D_{\min}(\alpha)=\min\{\alpha\bar{q},\bar{\alpha}q\}+\min\{\bar{\alpha}\bar{q},\alpha q\}$, which completes the proof. ∎
@@
$C_{BS-BC}=\bigcup_{\alpha\in[0,1/2]}\left\{\begin{array}[]{lc}(R_{1},R_{2},D)\in\mathbb{R}_{+}^{3}:&\\ R_{2}\leq 1-{\rm{H}}_{2}(\alpha\ast q)&\\ R_{1}+R_{2}\leq{\rm{H}}_{2}(\alpha)+1-{\rm{H}}_{2}(\alpha\ast q)&\\ D\geq\min\{\alpha\bar{q},\bar{\alpha}q\}+\min\{\bar{\alpha}\bar{q},\alpha q\}\end{array}\right\}.$ (23)
@@
and
$\sum_{i=1}^{n}\mathbb{E}[d(\hat{S}_{i},S_{i})]\leq nD.$
Using Fano’s inequality, we have that
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}H(X_{i}|\hat{X}^{n})\leq\frac{1}{n}\sum_{i=1}^{n}H(X_{i}|\hat{X}_{i})\leq\frac{1}{n}\sum_{i=1}^{n}H(p_{i})+p_{i}\log(|\mathcal{X}|-1)$
$\leq$ $H\bigg{(}\frac{1}{n}\sum_{i=1}^{n}p_{i}\bigg{)}=H\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[d(\hat{S}_{i},S_{i})]\bigg{)}.$ (24)
Hence we obtain
$\frac{1}{n}H(X^{n})-\frac{1}{n}I(X^{n};Z^{n})\leq H\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[d(\hat{S}_{i},S_{i})]\bigg{)}.$ (25)
Combining the fact that $H(X^{n})=nR$ with $\frac{1}{n}I(X^{n};Z^{n})\leq 1-H_{2}(q)$, we get $H_{2}^{-1}\big{(}R-1+H_{2}(q)\big{)}\leq D$. ∎
@@ IV-B Example 2
We consider a binary communication channel with a multiplicative Bernoulli state, where the signal received by the ComRx is given by $Y=SX$ with binary alphabets $\mathcal{X}=\mathcal{S}=\mathcal{Y}=\{0,1\}$ and the state $S$ is Bernoulli$(q)$, for $q\in(0,1)$. The signal received by the SenRx is expressed as the sum of two parts, where one part is the signal reflected by the ComRx and the other part is the directly radiated signal by of the ISAC Tx. The first part can be expressed as the product of a binary channel with a multiplicative Bernoulli state and an attenuation factor $a\in(0,1)$ and the second part can be expressed as a binary symmetric channel $\big{(}\mathrm{BSC}(e)\big{)}$ with error probability $e\in[0,1]$. Hence, the sensing channel can be expressed as $Z=aSX+X\oplus N$ with alphabets $\mathcal{Z}=\{0,1,a,a+1\}$, where $a\in(0,1)$, $N$ is Bernoulli$(e)$, and $e\in[0,1]$. We still consider the Hamming distortion measure $d(s,\hat{s})=s\oplus\hat{s}$ to characterize the sensing accuracy.
Applying the theoretical results in Section III to Example 2, we obtain the rate-distortion functions in Appendix E and the corresponding curves as shown in Fig. 5.
@@ V Conclusion
In this paper, we considered a bistatic ISAC system in which the sensing receiver is unaware of the sent message. Under this model, we provided a multi-letter characterization of the capacity-distortion function and transformed the resulting infinite-dimensional optimization problem into an infinite sequence of finite-dimensional optimization problems. For ease of calculation, we derive three single-letter lower bounds based on the three decoding-and-estimation (DnE) strategies: blind estimation, partial-decoding-based estimation, and full-decoding-based estimation, respectively. Additionally, three single-letter upper bounds are proposed by introducing a genie and leveraging Csiszar sum identity and the lossy source coding, respectively. In addition, we provided single-letter representations for the capacity-distortion functions in the two degraded cases, demonstrating the optimality of the partial-decoding-based estimation strategy achieved by superposition coding when the sensing channel is weaker than the communication channel. Furthermore, we extended these results to a bistatic ISAC degraded broadcast channel model, deriving the corresponding capacity-distortion region. Finally, we applied these theoretical results to two examples, illustrating the significance of the proposed bounds and the benefits of leveraging communications to assist sensing in a bistatic system.
@@ Appendix A Proof of Theorem 2
According to the definition of $C_{S,E}(D)$, we fix a stationary and ergodic stochastic process $X^{\prime}$ that achieves the $C_{S,E}(D)$, where the corresponding distortion satisfies $\lim_{k\rightarrow\infty}\frac{1}{k}\mathbb{E}\big{[}d\big{(}S^{k},\hat{S}^{*k}(Z^{k})\big{)}\big{]}\leq D$. Then, there must exist a $k$ such that $\frac{1}{k}\mathbb{E}\big{[}d\big{(}S^{k},\hat{S}^{*k}(Z^{k})\big{)}\big{]}\leq D-\epsilon$. Thus, we have $\frac{1}{k}I(X^{\prime k};Y^{\prime k}|S^{\prime k})\leq\sup_{p(x^{k})}\frac{1}{k}I(X^{k};Y^{k}|S^{k})$, and it implies $\lim_{k\rightarrow\infty}\frac{1}{k}I(X^{\prime k};Y^{\prime k}|S^{\prime k})\leq\sup\frac{1}{k}I(X^{k};Y^{k}|S^{k})$. Hence, according to the definition of $C(D-\epsilon)$ and Lemma 1, we get that $C_{S,E}(D)\leq C(D)$.
2) $C_{S,E}(D)>C(D)$
@@
$p(\tilde{X}_{j+1}^{m+1}=x_{j}^{m})=\frac{1}{k}\sum_{t=0}^{k-1}p\big{(}X_{j+1}^{m+1}(t)=x_{j}^{m}\big{)}=\frac{1}{k}\sum_{t=0}^{k-1}p\big{(}X_{j}^{m}(t)=x_{j}^{m}\big{)}=p(\tilde{X}_{j}^{m}=x_{j}^{m})$
@@
In order to more clearly express the connection between $\{\tilde{X}_{i}\}_{i=-\infty}^{\infty}$ and $\{X_{i}(t)\}_{i=-\infty}^{\infty}$, we introduce a random variable $T$ that is independent of everything else and uniformly distributed over $\{0,1,\dots,k-1\}$. Observe that $p\big{(}X_{i}(T)\big{)}=\sum_{t=0}^{k-1}p(T=t)p\big{(}X_{i}(T)|T=t\big{)}=\frac{1}{k}\sum_{t=0}^{k-1}p(X_{i}(t)\big{)}=\frac{1}{k}\sum_{t=0}^{k-1}p\big{(}X_{i+t}\big{)}$, for all $i$, thus, we can use $\{X_{i}(T)\}_{i=-\infty}^{\infty}$ instead of $\{\tilde{X}_{i}\}_{i=-\infty}^{\infty}$ in some of the proofs below. Similarly, we can define $p\big{(}S_{i}(T)\big{)}$ and $p\big{(}Y_{i}(T)\big{)}$.
@@
From (A) we can get, $\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}\Big{[}d\Big{(}S^{n}(j),\hat{S}^{*n}\big{(}Z^{n}(j)\big{)}\Big{)}\Big{]}\leq D$. Thus, for code length $n$, we obtain that the distortion satisfies $\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}\big{[}d\big{(}S^{n},\hat{S}^{*n}(Z^{n})\big{)}\big{]}\leq D$
In the end, we prove that $\lim_{n\rightarrow\infty}I(\tilde{X}^{n};\tilde{Y}^{n}|\tilde{S}^{n})\geq\frac{1}{k}I(X^{k};Y^{k}|S^{k})$. According to [46, Lemma 2, Lemma 3], we can get the above stationary construction way does not change the entropy limit, which means
@@
Furthermore, we get $C_{S,E}(D)>C(D)$ according to the definition of $C_{S,E}(D)$.

### v4

@@ III-D Capacity-distortion region for a bistatic ISAC broadcast channel model
In this subsection, we consider the bistatic ISAC broadcast systems, where the transmitter sends the message to the ComRx and the SenRx, the ComRx and the SenRx decode the information, and SenRx also needs to estimate the state of the channel. We propose a two-receiver state-dependent memoryless broadcast channel (SDMBC) model $\left(\mathcal{X},\mathcal{S},p\left(y,z|x,s\right),\mathcal{Y}\times\mathcal{Z}\right)$, which consists of four finite sets $\mathcal{X},\mathcal{S},\mathcal{Y},\mathcal{Z}$ and a collection of pmf $p\left(y,z|x,s\right)$ on $\mathcal{Y}\times\mathcal{Z}$. As shown in Fig. 3, the ISAC Tx sends a message $W=(W_{0},W_{1},W_{2})$, where $W_{0}$ is the common message to the ComRx and the SenRx, $W_{1}$ and $W_{2}$ are the private message to the ComRx and the SenRx, respectively.
@@ A $\left(2^{nR_{0}},2^{nR_{1}},2^{nR_{2}},n\right)$ code for SDMBC consists of
Since both capacity-distortion function and capacity-distortion region in this paper depend on the channel conditional pmf only through the conditional marginal pmfs, our capacity-distortion results for the physically degraded case defined by Markov chain are the same as those for stochastically degraded case.
Remark 7.
@@ IV-A Example 1
We consider a binary channel model, in which the signals received by the ComRx and the SenRx are given by $Z=Y=S\oplus X$ with binary alphabets $\mathcal{X}=\mathcal{S}=\mathcal{Y}=\{0,1\}$ and the state $S$ is Bernoulli$(q)$, for $q\in(0,1)$. The Hamming distortion measure $d(s,\hat{s})=s\oplus\hat{s}$ is employed to characterize the sensing accuracy.
In the following, we apply the theoretical results in section III to derive the capacity-distortion function. Let $\bar{e}=1-e$, $P(X=0)=p\in[0,1]$, and $e\ast p=e\bar{p}+\bar{e}p$. Besides, ${\rm{H}}_{2}(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$ is the binary entropy function and $\rm{H}_{2}^{-1}(\cdot)$ refers to the inverse function of the binary entropy function $\rm{H}_{2}(\cdot)$ with the variable taking the value in $[0,1/2]$. It is seen from the channel model in Example 1 that $Z$ is a degraded form of $Y$, i.e., the Markov chain $(X,S)-Y-Z$ holds. It follows from Theorem 6 that the capacity-distortion function is given by
$C(D)=\max_{P_{UX}}\big{\{}H(X|U)+I(U;Z)\big{|}\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(U,Z)\big{)}\big{]}\leq D\big{\}},$ (22)
@@
Without loss of generality, we assume that $q\geq 1/2$. Applying Theorem 3 to Example 1, we obtain that $R=I(X;Y|U,S)+I(U;Z)=H(X|U)+H(Z)-H(Z|U)$ and the distortion $\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(U,Z)\big{)}\big{]}=\sum_{u,z}\min_{s}p(s,u,z)$. Similar to the proof of the capacity region of a binary symmetric broadcast channel (BS-BC), we get that $R={\rm{H}}_{2}(\alpha)+1-{\rm{H}}_{2}(\alpha\ast q)$ is achievable by applying the superposition coding and successive cancellation decoding, i.e., $X=U\oplus V$, where $U$ and $V$ are independent, $U$ is Bernoulli$(\beta)$, and $V$ is Bernoulli$(\alpha)$ for $\beta\in[0,1]$. Correspondingly, the distortion $D=\min\{\alpha\bar{q},\bar{\alpha}q\}+\min\{\bar{\alpha}\bar{q},\alpha q\}$.
In the following, we prove the converse. By leveraging the fact that there exists an $\alpha\in[0,1/2]$ such that $0\leq H(Z)-H(Z|U)\leq 1-{\rm{H}}_{2}(q)=1-{\rm{H}}_{2}(\alpha\ast q)$ and the Mrs. Gerber’s Lemma in [45], we get that ${\rm{H}}_{2}(\alpha\ast q)=H(Z|U)\geq{\rm{H}}_{2}\big{(}{\rm{H}}_{2}^{-1}\big{(}{\rm{H}}_{2}(X|U)\big{)}\ast q\big{)}$. According to the monotonicity of binary entropy function and combining the facts that $\alpha\ast q\geq 1/2$ and $\alpha\leq e$ if $\alpha\ast q\geq e\ast q$, we have $H(X|U)\leq{\rm{H}}_{2}(\alpha)$ for $\alpha\in[0,1/2]$, which leads to $R\leq{\rm{H}}_{2}(\alpha)+1-{\rm{H}}_{2}(\alpha\ast q)$. Furthermore, since $\left|\mathcal{U}\right|\leq|\mathcal{X}|+1=3$, we set $P(U=i)=\beta_{i},i=0,1,2$, where $\sum_{i=0}^{2}\beta_{i}=1$. Thus, we have $D\geq D_{\min}(\alpha_{0},\alpha_{1},\alpha_{2})=\sum_{i=0}^{2}\beta_{i}\big{(}\min\{\alpha_{i}\bar{q},\bar{\alpha}_{i}q\}+\min\{\bar{\alpha}_{i}\bar{q},\alpha_{i}q\}\big{)}$, where $p(x=0|u=i)=\alpha_{i},i=0,1,2$. From the above expression about $D_{\min}$, we further get that $D_{\min}(\alpha_{0},\alpha_{1},\alpha_{2})\geq\min_{\alpha_{i},i=0,1,2}\big{\{}\min\{\alpha_{i}\bar{q},\bar{\alpha}_{i}q\}+\min\{\bar{\alpha}_{i}\bar{q},\alpha_{i}q\}\big{\}}$. Therefore, there exists an $\alpha\in[0,1/2]$ such that $D_{\min}(\alpha_{0},\alpha_{1},\alpha_{2})\geq D_{\min}(\alpha)=\min\{\alpha\bar{q},\bar{\alpha}q\}+\min\{\bar{\alpha}\bar{q},\alpha q\}$, which completes the proof. ∎
@@
$C_{BC}=\bigcup_{\alpha\in[0,1/2]}\left\{\begin{array}[]{lc}(R_{1},R_{2},D)\in\mathbb{R}_{+}^{3}:&\\ R_{2}\leq 1-{\rm{H}}_{2}(\alpha\ast q)&\\ R_{1}+R_{2}\leq{\rm{H}}_{2}(\alpha)+1-{\rm{H}}_{2}(\alpha\ast q)&\\ D\geq\min\{\alpha\bar{q},\bar{\alpha}q\}+\min\{\bar{\alpha}\bar{q},\alpha q\}\end{array}\right\}.$ (23)
@@
See Appendix E for details. ∎
@@ IV-B Example 2
We consider a binary communication channel with a multiplicative Bernoulli state, where the signal received by the ComRx is given by $Y=SX$ with binary alphabets $\mathcal{X}=\mathcal{S}=\mathcal{Y}=\{0,1\}$ and the state $S$ is Bernoulli$(q)$, for $q\in(0,1)$. The signal received by the SenRx is expressed as the sum of two parts, where one part is the signal reflected by the ComRx and the other part is the directly radiated signal by of the ISAC Tx. The first part is expressed as the product of a binary channel with a multiplicative Bernoulli state and an attenuation factor $a\in(0,1)$ and the second part is expressed as a binary symmetric channel $\big{(}\mathrm{BSC}(e)\big{)}$ with error probability $e\in[0,1]$. Hence, the sensing channel is expressed as $Z=aSX+(X\oplus N)$ with alphabets $\mathcal{Z}=\{0,1,a,a+1\}$, where $a\in(0,1)$, $N$ is Bernoulli$(e)$, and $e\in[0,1]$. We still consider the Hamming distortion measure $d(s,\hat{s})=s\oplus\hat{s}$ to characterize the sensing accuracy.
Applying the theoretical results in Section III to Example 2, we obtain the rate-distortion functions in Appendix F and the corresponding curves as shown in Fig. 5.
@@ V Conclusion
In this paper, we considered a bistatic ISAC system in which the sensing receiver is unaware of the sent message. Under this model, we provided a multi-letter characterization of the capacity-distortion function and transformed the resulting infinite-dimensional optimization problem into an infinite sequence of finite-dimensional optimization problems. For ease of calculation, we derive three single-letter lower bounds based on the three decoding-and-estimation (DnE) strategies: blind estimation, partial-decoding-based estimation, and full-decoding-based estimation, respectively. Additionally, three single-letter upper bounds are proposed by introducing the genie estimators and the lossy source coding. In particular, for the special case where the channel to SenRx is degraded with respect to the channel to ComRx, we provide the single-letter characterization of the capacity-distortion function, demonstrating the optimality of the partial-decoding-based estimation strategy achieved by superposition coding. Furthermore, we extended these results to a bistatic ISAC degraded broadcast channel model, deriving the corresponding capacity-distortion region. Finally, we applied these theoretical results to two examples, illustrating the significance of the proposed bounds and the benefits of leveraging communications to assist sensing in a bistatic system.
@@ Appendix A Proof of Theorem 2
Let a stationary and ergodic stochastic process $X^{\prime}$ achieves $C_{S,E}(D)-\epsilon$, where the corresponding distortion satisfies $\lim_{k\rightarrow\infty}\frac{1}{k}\mathbb{E}\big{[}d\big{(}S^{k},\hat{S}^{*k}(Z^{k})\big{)}\big{]}=D$ according to the definition of $C_{S,E}(D)$ and Lemma 1. Then, given $\epsilon>0$, for sufficiently large $k$, $\frac{1}{k}\mathbb{E}\big{[}d\big{(}S^{k},\hat{S}^{*k}(Z^{k})\big{)}\big{]}\leq D+\epsilon$. Furthermore, since $I(X^{\prime};Y^{\prime}|S^{\prime})+\epsilon=\lim_{k\rightarrow\infty}\frac{1}{k}I(X^{k};Y^{k}|S^{k}),$ there exists a $p(x^{k})$ such that $\frac{1}{k}I(X^{k};Y^{k}|S^{k})\geq I(X^{\prime};Y^{\prime}|S^{\prime})-\epsilon$ for sufficiently large $k$. Therefore, we conclude that $C(D+\epsilon)=\lim_{k\to\infty}\sup_{p(x^{k})}\frac{1}{k}I(X^{k};Y^{k}|S^{k})\geq I(X^{\prime};Y^{\prime}|S^{\prime})-\epsilon=C_{S,E}(D)-\epsilon$. Taking $\epsilon\to 0$ and using Lemma 1, we get $C(D)\geq C_{S,E}(D).$
2) $C_{S,E}(D)\geq C(D)$
@@
$p(\tilde{X}_{j+1}^{m+1}=x_{j}^{m})=\frac{1}{k}\sum_{t=0}^{k-1}p\big{(}X_{j+1}^{m+1}(t)=x_{j}^{m}\big{)}$
$=$ $\frac{1}{k}\sum_{t=0}^{k-1}p\big{(}X_{j}^{m}(t)=x_{j}^{m}\big{)}=p(\tilde{X}_{j}^{m}=x_{j}^{m})$
@@
In order to more clearly express the connection between $\{\tilde{X}_{i}\}_{i=-\infty}^{\infty}$ and $\{X_{i}(t)\}_{i=-\infty}^{\infty}$, we introduce a random variable $T$ that is independent of everything else and uniformly distributed over $\{0,1,\dots,k-1\}$. Observe that
$p\big{(}X_{i}(T)\big{)}=\sum_{t=0}^{k-1}p(T=t)p\big{(}X_{i}(T)|T=t\big{)}=\frac{1}{k}\sum_{t=0}^{k-1}p(X_{i}(t)\big{)}=\frac{1}{k}\sum_{t=0}^{k-1}p\big{(}X_{i+t}\big{)},$
for all $i$, thus, we use $\{X_{i}(T)\}_{i=-\infty}^{\infty}$ instead of $\{\tilde{X}_{i}\}_{i=-\infty}^{\infty}$ in the proof below. Similarly, we can define $p\big{(}S_{i}(T)\big{)}$ and $p\big{(}Y_{i}(T)\big{)}$.
@@
From (A) we get that $\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}\Big{[}d\Big{(}S^{n}(j),\hat{S}^{*n}\big{(}Z^{n}(j)\big{)}\Big{)}\Big{]}\leq D$. Thus, for code length $n$, we obtain that the distortion satisfies $\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}\big{[}d\big{(}S^{n},\hat{S}^{*n}(Z^{n})\big{)}\big{]}\leq D$
In the end, we prove that $\lim_{n\rightarrow\infty}I(\tilde{X}^{n};\tilde{Y}^{n}|\tilde{S}^{n})\geq\frac{1}{k}I(X^{k};Y^{k}|S^{k})$. According to [46, Lemma 2, Lemma 3], we get the above stationary construction way does not change the entropy limit, which means
@@
Furthermore, we get $C_{S,E}(D)\geq C(D)$ according to the definitions of $C_{S,E}(D)$ and $C(D)$.